OpenStudy (anonymous):

give an algebraic equation in which no coefficients or constant term is \[\pi\] but it has a root equal to \[\pi\]

OpenStudy (anonymous):

any one here???

OpenStudy (anonymous):

dx - c = 0 where d is the diameter c i s the circumference of any circle

OpenStudy (anonymous):

what say?

OpenStudy (anonymous):

circumference itself contains pi

OpenStudy (anonymous):

but good try

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

\[ds=rd \theta\]\[s=2 \pi r\]

OpenStudy (anonymous):

are u denying this???

OpenStudy (anonymous):

the value of pi is derived from circumference

OpenStudy (anonymous):

do u agree the basic formula that i've written ??

OpenStudy (anonymous):

pi is DEFINED as the ratio of circumference to the diameter.....that is how u derive the formula for circumference

OpenStudy (anonymous):

obviously not

OpenStudy (anonymous):

the people who made this formula did not use line integrals to calculate length..instead they defined the ratio of the circumference to the diameter as pi

OpenStudy (anonymous):

the state ment u gave is for the class 9th or 10th standard

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

it is historical

OpenStudy (anonymous):

but it has no meaning..

OpenStudy (anonymous):

then u give the answer

OpenStudy (anonymous):

i need an equation like f(x)=c where all coefficients and constants must be rational no.

OpenStudy (anonymous):

where did u mention rational nos?

OpenStudy (anonymous):

now i'm mentioning it.

OpenStudy (anonymous):

see @himanish u are a very good student, tell me correctly do u think that what u have answered is acceptable??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

bt thats only when i didnt know abt the rational no thing

OpenStudy (anonymous):

i don't think so. if the ans could be so easy then i would not post it here.

OpenStudy (anonymous):

bhalo

OpenStudy (anonymous):

think some different approach i think u can do it.

OpenStudy (anonymous):

\[\pi = 3\sum_{n=0}^{\infty}[\left(\begin{matrix}2n \\ n\end{matrix}\right) / 16^{n}(2n+1)]\]

OpenStudy (anonymous):

what to do?

OpenStudy (anonymous):

this is pi

OpenStudy (anonymous):

well but how can u write algebraic equation??

OpenStudy (anonymous):

[ix - ln(-1)] = 0

OpenStudy (anonymous):

the above series is algebraic

OpenStudy (anonymous):

thats great..

OpenStudy (anonymous):

not the second one

OpenStudy (anonymous):

[ix - ln(-1)] = 0 is this right?

OpenStudy (anonymous):

yes but it contains i

OpenStudy (anonymous):

dont u use i in algebra?

OpenStudy (anonymous):

u used \[e^{i \pi}=-1\]

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

is i rational???

OpenStudy (anonymous):

[x - |ln(-1)| ] = 0 now?

OpenStudy (anonymous):

|ln (-1) | is pi isnt it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is it okay?

OpenStudy (anonymous):

so which one is better the first or it??

OpenStudy (anonymous):

this one is shorter

OpenStudy (anonymous):

first one is wrong.

OpenStudy (anonymous):

the series?

OpenStudy (anonymous):

again here is some mistakes but i can ignore it

OpenStudy (anonymous):

so the answer?

OpenStudy (anonymous):

no dx-c=0 is wrong

OpenStudy (anonymous):

x=|ln(-1)| x=(2n+1)pi x=pi when n=1

OpenStudy (anonymous):

what was the answer u were lookin for?

OpenStudy (dumbcow):

its impossible there is no polynomial function with a transcendental root

OpenStudy (anonymous):

there are a lot of answers.

OpenStudy (anonymous):

like?

OpenStudy (anonymous):

\[1-x^2/2!+x^4/4!-....=1\]

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Transcendental_numbers pi cannot be a root to a function with all rational coefficients

OpenStudy (anonymous):

what do u think @dumbcow about the above equation??

OpenStudy (anonymous):

dipankar how does it come to pi?

OpenStudy (anonymous):

put pi over there u will find RHS=LHS

OpenStudy (anonymous):

THAT IS VALID FOR ANY X

OpenStudy (dumbcow):

you have an infinite series, a algebraic equation is defined as finite number of terms

OpenStudy (anonymous):

sorry right hand side will be -1

OpenStudy (anonymous):

this answer is invalid

OpenStudy (anonymous):

why invalid??

OpenStudy (dumbcow):

you can always write pi as an infinite series

OpenStudy (anonymous):

like i said above

OpenStudy (anonymous):

6arcsin(0.5)

OpenStudy (anonymous):

sin is not algebraic.

OpenStudy (anonymous):

bt its expressible as an infinite algebraic series

OpenStudy (anonymous):

any function can be expressed as an algebraic series

OpenStudy (anonymous):

by taylor series

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

so what about |ln(-1)|??

OpenStudy (anonymous):

it can be expanded 2

OpenStudy (anonymous):

its constant.lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pi